# Instance-level quantitative saliency in multiple sclerosis lesion segmentation

### v2

@@ 2.5 Experiments
To understand the model’s level of specificity in segmenting MS lesions in the WM, we designed three additional qualitative experiments on a batch of 10 patients, in which we observed the model’s behavior on synthetic lesions (results in Section 3.4). A first experiment was conducted, moving a clearly visible WM lesion to a different part of the WM, which originally presented no lesions. A similar approach was followed when inserting the same lesion outside the skull. A third experiment consisted of a compromise: the lesion, along with part of its surrounding healthy tissue ($3mm$ from lesion border), was moved outside the skull. For all these cases, the model’s prediction and saliency maps were examined.
@@
In light of the results observed during the described tests, we designed a more specific experiment on the entire test set: the analysis of the needed amount of contextual information surrounding a lesion to obtain a segmentation (results in Section 3.5). We selected lesions with size close to the average of the entire dataset, i.e., from $90$ to $120mm^{3}$, obtaining a total of 178 lesions. Initially, all voxel intensities were set to zero, but those of the lesion: this step was called iteration 0. Then, we gradually reassigned the original intensity to surrounding voxels through morphological 3D dilation, iterating this process 25 times. With the final iteration, the model was seeing a volume of surrounding tissue at a maximum of $25mm$ distance from the lesion’s edges. At each iteration, the following metrics were recorded:
  • 1.
@@
3 Results
3.1 Maximum versus average aggregation for quantitative saliency maps
@@
Fig. 7: Saliency maps based on SG generated for an extensive lesion, following the average (a) and the maximum (b) methodology. Saliency maps generated for a smaller lesion, following the average (c) and the maximum (d) methodology. The thumbnails illustrate the lesion domains $\Omega$ used during the computation.
3.2 Assessing the contributions of input MR sequences using gradient-based saliency maps
@@
Fig. 8: Example of a FLAIR image (a), the lesion domain $\Omega$ (b), and the corresponding saliency map obtained with the proposed adaptation of SmoothGrad isolating positive (c) and negative (d) gradients. Values in [-0.05, 0.2] are not displayed to focus on most significant saliency.
Fig. 9: Example of an MPRAGE image (a), the lesion domain $\Omega$ (b), and the corresponding saliency map obtained with the proposed adaptation of SmoothGrad isolating positive (c) and negative (d) gradients. Values in [-0.1, 0.1] are not displayed to focus on the most significant saliency.
Fig. 10: Saliency maps obtained with the proposed adaptation of SmoothGrad for two close lesions (a) and (b). Heatmaps obtained with the proposed adaptation of Grad-CAM++ for the same two lesions (c) and (d). The thumbnails illustrate the lesion domains $\Omega$ used during the computation.
3.3 Statistical distribution of gradient values for TPs, FPs, FNs and TNs
Fig. 11 reports the distribution of maximum (a) and minimum (b) gradients values — with respect to FLAIR — for TN, FN, FP and TP volumes. The median value and 95% CI for positive values in each group of volumes were, respectively: $0.465\pm 0.025$ (TN), $1.202\pm 0.036$ (FN), $1.630\pm 0.019$ (FP), and $1.997\pm 0.016$ (TP). The median value and 95% CI for negative values were, respectively: $-0.245\pm 0.008$ (TN), $-0.359\pm 0.010$ (FN), $-0.522\pm 0.008$ (FP), and $-0.639\pm 0.009$ (TP). The $p$-value of Mann Whitney U tests run on all pairs of groups was lower than $6\times 10^{-65}$.
Fig. 11: Violin plots representing the distribution of saliency maps maximum (a) and minimum (b) values. The four distributions refer to TN, FN, FP and TP volumes.
3.4 Sanity checks
@@
Fig. 12: The prediction score before the Softmax for the case of a lesion artificially placed in the WM (a), in the background without (c) and with (e) a part of the surrounding WM. The corresponding saliency maps (b), (d) and (f), respectively. Only the synthetic lesion in the WM triggered a detection.
3.5 Experiment on contextual information
@@
The experiment on contextual information showed that, compared to a lesion without any surroundings (i.e., constant background values outside the lesion mask), the prediction score for a lesion increased when including voxels belonging to its perilesional healthy tissue. The prediction score reached a plateau when adding tissue distant $12-15mm$ from the lesion border. A minimum of $7mm$ of healthy perilesional tissue allowed to correctly segment all the 178 WM lesions, as reported in Fig. 13. The standard deviation reached a peak at $3mm$ distance from lesion’s border, meaning that some lesions already presented high scores while others had not yet been segmented.
@@
Fig. 13: (A): FLAIR masked out with dilation steps 1, 5 and 24 (top), and the corresponding output probability maps (bottom). (B): Plots representing the number of segmented lesions (top) and the average and standard deviation across patients of the mean prediction score (bottom) at each dilation step.
4 Discussion
Our study demonstrated the successful adaptation of two existing XAI methods to the task of semantic segmentation. Specifically, we are able to generate quantitative explanation maps for single segmented instances, such as MS lesions.
@@ 4 Discussion
The distribution of values in the saliency maps generated with SG showed that FLAIR imaging had a more significant contribution to the segmentation of lesions, compared to MPRAGE. This resonates with the fact that FLAIR offers a better contrast for common MS lesions in the WM, such as periventricular and sub-cortical, since the signal from the cerebrospinal fluid is suppressed. Only a part of these lesions are clearly visible in T1-weighted images, such as MPRAGE, which are preferred to detect cortical plaques [33, 22].
@@
Both of our XAI methods showed that voxels far from a lesion domain $\Omega$ did not appear to impact the prediction of voxels belonging to $\Omega$ (Fig. 10). Since we employed a CNN, each network unit did not depend on the entire input, but on a region called receptive field. [17] described that the effective area of a receptive field (effective receptive field, or ERF) starts as small, and then grows during training. Furthermore, the same study described skip-connections, part of a U-Net, as a cause of the reduction of receptive field’s size. In our particular case, one of the possible interpretations could be that useful features to segment a lesion were close to the lesion itself and its neighborhood, so the learned receptive field is small. For instance, it is likely that re-training our model by including the images of the contextual experiment (e.g., until dilation step 25) and labeling them as background, would lead the model to have a larger ERF. Saliency maps, after such re-training, may show a higher attention to neighbouring tissue at a greater distance from a lesion border.
@@
However, the last experiment on contextual information revealed that high and stable prediction scores were related to the amount of contextual healthy brain tissue from the perilesional volume (Fig. 13). Enriching the context around the lesion in the input resulted in an increase of prediction scores up to a distance of $12-15mm$ from the lesion border. Past this distance, additional voxels no longer impact the prediction as seen by the plateau in Fig. 13(B). A possible conclusion is that a patch size of $96mm^{3}$ could have been unnecessarily large for most lesions. Furthermore, it would be interesting to test if the patch size during training would have a potential influence on the plateau’s position of Fig. 13(B). We might expect that bigger patches would make the model require more contextual information to segment a lesion.
5 Conclusion
We adapted two XAI methods to provide quantitative instance-level explanations for segmentation, which we applied to the specific case of MS lesion segmentation. The analysis of the explanation maps and additional tests revealed fundamental insights into the decision mechanism of a deep neural network, which could also be used to orient architecture choices (e.g., patch size), and to boost the sensitivity/specificity of the model. The acquired new knowledge can be an important step in facilitating AI integration into clinical practice. The proposed methods can potentially be applied to various segmentation tasks outside the medical imaging field, such as autonomous driving and robotics.
Acknowledgments

### v3

@@ Methods
To understand the U-Net’s level of specificity in segmenting MS lesions in the WM, we designed three additional qualitative experiments on a batch of 10 patients, in which we observed the U-Net’s behavior on synthetic lesions (results in Section Sanity checks). A first experiment was conducted, moving a clearly visible WM lesion to a different part of the WM, which originally presented no lesions. A similar approach was followed when inserting the same lesion outside the skull. A third experiment consisted of a compromise: the lesion, along with part of its surrounding healthy tissue ($3mm$ from lesion border), was moved outside the skull. For all these cases, the U-Net’s prediction and saliency maps were examined.
@@
In light of the results observed during the described tests, we designed a more specific experiment on the entire test set, using all the three trained networks: the analysis of the needed amount of contextual information surrounding a lesion to obtain a segmentation (results in Section Experiment on contextual information). We selected lesions with size close to the average of the entire dataset, i.e. from $90$ to $120mm^{3}$, obtaining a total of 191 (U-Net), 173 (nnU-Net), and 161 (Swin UNETR) TP lesions. Initially, all voxel intensities were set to zero, but those of the lesion: this step was called iteration 0. Then, we gradually reassigned the original intensity to surrounding voxels through morphological 3D dilation, iterating this process 35 times. With the final iteration, the models were seeing a volume of surrounding tissue at a maximum of $35mm$ distance from the lesion’s edges. At each iteration, the following metrics were recorded:
@@
Results
The trained lesion segmentation models —U-Net, nnU-Net, Swin UNETR— achieved, respectively, a test Dice score of 0.60, 0.62, 0.66, a normalized Dice score of 0.71, 0.78, 0.80, a true positive rate of 79%, 78%, and 85%, a false discovery rate of 37%, 38%, and 36%, a false negative rate of 20%, 22%, and 14%. The count of TP, FP and FN examples was 3050, 1818, 789 for the U-Net, 3112, 1954, 880 for the nnU-Net, and 3516, 2018, 598 for the Swin UNETR.
Maximum versus average aggregation for quantitative saliency maps
@@ Results
Figure 7: Saliency maps (computed for U-Net) based on SG generated for an extensive lesion, following the average (a) and the maximum (b) methodology. Saliency maps generated for a smaller lesion, following the average (c) and the maximum (d) methodology. The gradients intensity difference between (a,c) and (b,d) can be observed from the color scales range next to each map, increasing approximately from (-0.1, 0.3) to (-1, 2.5). The thumbnails illustrate the lesion domains $\Omega$ used during the computation.
Assessing the contributions of input MR sequences using gradient-based saliency maps
@@
Figure 8: Example of a FLAIR image (a), the lesion domain $\Omega$ (b), and the corresponding saliency map (computed for U-Net) obtained with the proposed adaptation of SmoothGrad isolating positive (c) and negative (d) gradients. Values in [-0.05, 0.2] are not displayed to focus on most significant saliency.
Figure 9: Example of an MPRAGE image (a), the lesion domain $\Omega$ (b), and the corresponding saliency map (computed for U-Net) obtained with the proposed adaptation of SmoothGrad isolating positive (c) and negative (d) gradients. Values in [-0.1, 0.1] are not displayed to focus on the most significant saliency.
Figure 10: Saliency maps (computed for U-Net) obtained with the proposed adaptation of SmoothGrad for two close lesions (a) and (b). Heatmaps obtained with the proposed adaptation of Grad-CAM++ for the same two lesions (c) and (d). The thumbnails illustrate the lesion domains $\Omega$ used during the computation.
Statistical distribution of gradient values for TPs, FPs, FNs and TNs
Fig. 11 reports the distribution of maximum (a) and minimum (b) gradients values — with respect to FLAIR — for TN, FN, FP and TP volumes. The median value and 95% CI for positive values in each group of volumes were, respectively: $0.465\pm 0.025$ (TN), $1.202\pm 0.036$ (FN), $1.630\pm 0.019$ (FP), and $1.997\pm 0.016$ (TP). The median value and 95% CI for negative values were, respectively: $0.245\pm 0.008$ (TN), $0.359\pm 0.010$ (FN), $0.522\pm 0.008$ (FP), and $0.639\pm 0.009$ (TP). The Mann Whitney U tests run on all pairs of groups reported a $p$-value < 0.001.
Figure 11: Violin plots representing the distribution of saliency maps (computed for U-Net) maximum (a) and minimum (b) values. The four distributions refer to TN, FN, FP and TP volumes.
Sanity checks
The saliency map generated from a volume (in the WM) with no lesions resulted in gradient values in the range of TN examples (Fig. 11), that is about 5 times smaller than the average one obtained in TP cases. An example of this finding is illustrated in Fig. 12.
Figure 12: A region in FLAIR with healthy white matter (a) and the corresponding instance-level saliency map (computed for U-Net) obtained with the proposed adaptation of SmoothGrad (b). The color scale range is (-0.1, 0.3), well below the one in Figures 7b and 7d. The thumbnail illustrates the domain $\Omega$ used during the computation.
The case of a FLAIR with a single voxel domain $\Omega$, is shown in Fig. 13. Saliency values suggest the prediction of $\Omega$ is attributed to input voxels which are in the vicinity, and within the lesion. It appears clear that lesion voxels, in the input, which are farther from $\Omega$ present a lower contribution (as it was noticed when using the average aggregation method). However, not only the input voxel corresponding to $\Omega$ has an influence.
Figure 13: Axial view of an MS lesion, highlighting in red its center of mass used as domain $\Omega$ (a), and the corresponding voxel-level saliency map (computed for U-Net) obtained with the proposed adaptation of SmoothGrad (b).
@@
Figure 14: The prediction score before the Softmax for the case of a lesion artificially placed in the WM (a), in the background without (c) and with (e) a part of the surrounding WM. The corresponding saliency maps (b), (d) and (f), respectively. Only the synthetic lesion in the WM triggered a detection.
Experiment on contextual information
@@
The experiment on contextual information showed that, compared to a lesion without any surroundings (i.e. constant background values outside the lesion mask), the prediction score for a lesion increased when including voxels belonging to its perilesional healthy tissue. For U-Net and nnU-Net, the prediction score reached a similar plateau when adding tissue distant 1215mm12-15mm from the lesion border. A minimum of $7mm$ of healthy perilesional tissue allowed to correctly detect all the TP lesions, as reported in Fig. 15. The standard deviation reached a peak at $3mm$ distance from lesion’s border, meaning that some lesions already presented high scores while others had not yet been segmented. For Swin UNETR, the prediction score reached a plateau past $15mm$ from the lesion border, presenting a ripple around $6mm$. All the TP lesions were detected only when at least $13mm$ of healthy tissue were seen by the network.
@@
Figure 15: (A): FLAIR masked out with dilation steps 1, 5 and 24 (top), and the corresponding output probability maps computed for U-Net (bottom). (B): Plots representing the number of segmented lesions (left) and the average across patients of the mean prediction score (right) at each dilation step for three tested networks (the transparency for each line represents its standard deviation).
Discussion
We introduced novel methods to generate quantitative instance-level saliency maps, pushing the explanation capabilities of basic classification and segmentation saliency methods (SG and GradCAM++) to single instance segmentation. In medical image analysis, the latter is crucial to reveal the information used by a model to detect and segment a specific lesion of interest among other lesions of the same class, which is a very important task in daily radiology to assess disease development and response to treatment for all multi-lesional diseases (e.g., MS, metastatic cancer). In addition to providing explanations for a models’ decision, the proposed quantitative maps can be used to optimize a network’s architecture and to boost a semantic segmentation models’ performance. This is shown in Spagnolo et al. [45], where features from saliency maps were used to increase detection F1 score and positive predictive value.
Although the test Dice scores of U-Net and nnU-Net are below that of Swin UNETR and of other state of the art methods [55, 8], there are some aspects of our approach one should consider: (i) the target metric for the loss was the normalized Dice score, which was 0.71 at test (ii) the imaging data do not contain synthetic lesions, and are not skull stripped (iii) possibly, the volume of false positive predictions is relatively small and (iv) the computational cost of saliency maps for Swin UNETR is higher. The false positive predictions of the model have been visually examined, and are mostly located in cerebral sulci and gyri. The use of brain extraction tools on our data would also have contributed to the removal of some false positives, since a few of them were hyperintensities located in non-cerebral tissue.
@@ Discussion
The distribution of values in the saliency maps generated with SG showed that FLAIR imaging had a more significant contribution to the segmentation of lesions, compared to MPRAGE. This finding allowed us to check whether the model’s behavior is coherent with clinical practice. Indeed, it resonates with the fact that FLAIR offers a better contrast for common MS lesions in the WM, such as periventricular and sub-cortical, since the signal from the cerebrospinal fluid is suppressed. Only a part of these lesions are clearly visible in T1-weighted images, such as MPRAGE, which are preferred to detect cortical plaques [48, 34].
@@
Both of our XAI methods showed that voxels far from a lesion domain $\Omega$ did not appear to impact the prediction of voxels belonging to $\Omega$ (Fig. 10). Since we employed a CNN, each network unit did not depend on the entire input, but on a region called receptive field. Luo et al. [29] described that the effective area of a receptive field (effective receptive field, or ERF) starts as small, and then grows during training. Furthermore, the same study described skip-connections, part of a U-Net, as a cause of the reduction of receptive field’s size. In our particular case, one of the possible interpretations could be that useful features to segment a lesion were close to the lesion itself and its neighborhood, so the learned receptive field is small. For instance, it is likely that re-training our model by including the images of the contextual experiment (e.g., until dilation step 25) and labeling them as background, would lead the model to have a larger ERF. Saliency maps, after such re-training, may show a higher attention to neighbouring tissue at a greater distance from a lesion border.
@@
However, the last experiment on contextual information revealed that high and stable prediction scores were related to the amount of contextual healthy brain tissue from the perilesional volume (Fig. 15). For U-Net and nnU-Net, enriching the context around the lesion in the input resulted in an increase of prediction scores up to a distance of 1215mm12-15mm from the lesion border. Past this distance, additional voxels no longer impact the prediction as seen by the plateau in Fig. 15(B). A possible conclusion is that a patch size of $96mm^{3}$ could have been unnecessarily large for most lesions. Furthermore, it would be interesting to test if the patch size during training would have a potential influence on the plateau’s position of Fig. 15(B). We might expect that bigger patches would make the model require more contextual information to segment a lesion. The trajectory of prediction scores from Swin UNETR followed the same overall trend as the other two networks. However, certain architectural characteristics — such as the use of partitioned tiles ($8mm$ wide) applied to input patches — may make the model more sensitive to our experimental setup.
Regarding the potential of using the proposed XAI to improve model performance, the maximum and minimum values of XAI maps in Fig. 11 alone did not show enough discriminatory power between the different groups. However, a simple linear model relying on saliency radiomics allowed to reduce FPs, which is demonstrated in Spagnolo et al. (2025)[45] with a clear example of how our method can be relevant to the development and optimization of a segmentation model, and ultimately to the clinics.
Conclusion
We proposed novel XAI methods to provide quantitative instance-level explanations for segmentation, which we applied to the specific case of MS lesion segmentation. The analysis of the explanation maps and additional tests revealed fundamental insights into the decision mechanism of a deep neural network. The explanation maps were useful to understand the importance of the perilesional volume and improve the network’s classification performances. The following experiment on the contextual information exploited by the network can guide architecture choices, such as patch size. The acquired new knowledge is crucial for AI engineers and clinical researchers, which constitutes an important step in facilitating AI integration into clinical practice. The proposed methods can potentially be applied to various segmentation architectures and tasks outside the medical imaging field, such as autonomous driving and robotics. Future research could leverage uncertainty maps to generate XAI maps of false negative examples, and further boost performance. Additionally, the pruning of false predictions could be studied separately for different areas of the brain.
Data Availability
The datasets generated during and/or analysed during the current study are available from the corresponding author on reasonable request.
Acknowledgements
@@
Author contributions statement
F.S., V.A. and A.D. conceived the experiments, F.S. conducted the experiments, L.M-G, M.O.P. and F.S. processed and organized the data, F.S. and N.M. created the software used in the work. V.A., A.D. and C.G. supervised the work. F.S. led the writing of the manuscript together with V.A. and A.D. All authors reviewed the manuscript.
Additional information
Competing interests
The University Hospital Basel (USB) and the Research Center for Clinical neuroimmunology and Neuroscience (RC2NB), as the employers of Cristina Granziera, have received the following fees which were used exclusively for ( research support from Siemens, GeNeuro, Genzyme-Sanofi, Biogen, Roche. They also have received advisory board and consultancy fees from Actelion, Genzyme-Sanofi, Novartis, GeNeuro, Merck, Biogen and Roche; as well as speaker fees from Genzyme-Sanofi, Novartis, GeNeuro, Merck, Biogen and Roche.